OpenStudy (anonymous):

f(x)=3x-2 find f(1/x)

OpenStudy (anonymous):

This problem in words is asking you what the inverse of (3x-2) is. :-) hint: Inverse of 4 is \(\large\frac{1}{4}\) or 0.25

OpenStudy (anonymous):

1/3x-2??

OpenStudy (anonymous):

:-O Parenthesis are important! ;-)

OpenStudy (anonymous):

(1/3x)-2

OpenStudy (anonymous):

\[1/3x-2 \neq 1/(3x-2) = \frac{1}{3x-2}\]

OpenStudy (anonymous):

\[1/3x-2 = \frac{1}{3x}-2\]

OpenStudy (anonymous):

but my text gives a different answer.

OpenStudy (anonymous):

(3/x)-2

OpenStudy (anonymous):

Ah I didn't understand what you originally meant then, \[3(\frac{1}{x})-2\]

OpenStudy (anonymous):

You're are putting the inverse of x into the input, gotcha

OpenStudy (anonymous):

it's not asking for the inverse

OpenStudy (anonymous):

o wait you figured it ou nvm lol

OpenStudy (anonymous):

@Outkast3r09 Yeah I see that now.

OpenStudy (anonymous):

I do not understand?

OpenStudy (anonymous):

3 * 1/x is the same thing as 3 divided by x Numerical proof: (pretending x is 2) 3 * (1/2) = 1.5 3 / 2 = 1.5

OpenStudy (anonymous):

Or if you want to take it up a notch, lets say x = 0.5, which is 1/2 3 * (1/(1/2)) = 3 * (1/(0.5)) = 3 * 2 = 6

OpenStudy (anonymous):

Now do you understand @Hollywood_chrissy ?

OpenStudy (anonymous):

Fact: Multiplying by a reciprocal of a number is the same thing as dividing by that number \[\frac{3}{1}\cdot \frac{1}{x} = \frac{3}{x}\]

OpenStudy (anonymous):

a little.

OpenStudy (anonymous):

3 * 1 = 3 1 * x = x

OpenStudy (anonymous):

and then subtract 2

OpenStudy (anonymous):

\(\huge\frac{3}{x} - 2\)

OpenStudy (anonymous):

thanks sale

OpenStudy (anonymous):

np, sry about the confusion at the start